OpenStudy (anonymous):

Can someone confirm my calculus answer regarding integrals? y'+(2x-1)y=0 I got y=e^(-x^2+x)

OpenStudy (anonymous):

gotcha. I'm running into a problem though, it's supposed to satisfy a "boundary condition" (not entirely sure what that is) of y(1)=2, but when I plug in x=1 I get y=1

OpenStudy (loser66):

I think we missed C but C is in the front of e I mean \(y = Ce^{-x^2+x}\) so, if y (1) =2, just plug x =1, y =2 into the solution \[2=Ce^{-1^2+1}=Ce^0= C\] therefore C =2 Now, plug it back to the solution \[y = Ce^{-x^2+x}= 2e^{-x^2+x}\]

OpenStudy (anonymous):

why is it C*e instead of C+e

OpenStudy (anonymous):

also if we fill in x with 1, wouldn't we get Ce^2?

OpenStudy (anonymous):

also can you explain how to set this up? http://puu.sh/9gqir/f2a80deb8e.png

OpenStudy (loser66):

ok, y'= -(2x-1)y -->\(\dfrac{dy}{dx}=-(2x-1)y\rightarrow \dfrac{dy}{y}=(-2x +1)dx\) Integral both sides \[lny = -x^2+x +C\] take e both sides \[\Large y = e^{the~ whole~ thing}= e^{-x^2+x+C}= e^C*e^{-x^2+x}\] e^C = C because it is a constant so that C * e^(-x^2+x} not + as usual

OpenStudy (loser66):

for the previous one, re read my comment \[e^{-x^2+x}= e^0\] the minus sign is not include in x^2 --> -x^2 = -1^2 , then +1 =0

OpenStudy (loser66):

for the next problem , you need to write out the equation of the distance first, then take derivative of that function to get the velocity function. That's the way you make up ODE

OpenStudy (anonymous):

A little late. Refer to a solution using Mathematica 9.